OpenStudy (tiniloisx):

Please help solving a problem, don't give me the answer. Work me through it! Medal + Fan! :] Julia made $130 for 6 hours of work. At the same rate, how many hours would she have to work to make $399.

OpenStudy (tyrenhere):

What you have to do is find out how much money she makes in a certain amount of time. How can you do that?

OpenStudy (tiniloisx):

Please if I answer wrongly (most likely) keep in mind I suck at math and am pretty dumb in this subject but maybe divide the number of hours by the amount she made in that time? so 130/6?

OpenStudy (tyrenhere):

That's correct! :3

jhonyy9 (jhonyy9):

Tyrenhere what mean this 3 ?

OpenStudy (tiniloisx):

Omg okay so I got 21.6 for 130/6.

OpenStudy (tyrenhere):

It's like a face. the : are eyes and the 3 are lips

OpenStudy (tyrenhere):

And so that means she make $21.67 an hour.

OpenStudy (tyrenhere):

We wanna know how much she has to work to get $399 dollars so what do you think we should do now?

jhonyy9 (jhonyy9):

TyrenHere good idea - continue it please

OpenStudy (tyrenhere):

What's a good idea?

OpenStudy (tyrenhere):

@jhonyy9

jhonyy9 (jhonyy9):

this And so that means she make $21.67 an hour.

OpenStudy (tyrenhere):

Oh

jhonyy9 (jhonyy9):

that you calculi how many get for one hour

OpenStudy (tyrenhere):

Yeah, She specifically asked for me to walk her through it so I'm waiting on he reply xD

jhonyy9 (jhonyy9):

tini please collaborate - bc. here on OS not is possible give you just direct answer

OpenStudy (tiniloisx):

Sorry had to use the bathroom lol

OpenStudy (tyrenhere):

I would have gave her the direction answer to be honest xD but she asked me to explain but what's the next step now that we know how much is made an hour?

jhonyy9 (jhonyy9):

ok. just try understanding what wrote above TyrenHere

OpenStudy (tiniloisx):

Are we multiplying now?

jhonyy9 (jhonyy9):

so now that you know how many she get for one hour so you know she how many get for how many minutes ?

jhonyy9 (jhonyy9):

just bc. 399 not is divisibly bye 130 exactly so you need to know how many she get for minutes

OpenStudy (tyrenhere):

Jhonyy, you can take it from here ;) If it doesn't work out just message me for help! :3

jhonyy9 (jhonyy9):

the easy part that she get for 6 hours 130 us dollars so multiplied by 3 she get 390 dollars in 3*6 hours - yes ?

OpenStudy (tyrenhere):

You shoud be able to voice chat when answer question. This would go so much quicker.

OpenStudy (tyrenhere):

But you know there are people with dirty mouths so nevermind LOL

jhonyy9 (jhonyy9):

TyrenHere - do you can ending it now please explained it step by step how she get the needed 9 dollars ?

OpenStudy (tyrenhere):

I'm kind of preoccupied I'm doing my own math quiz ATM.

OpenStudy (tyrenhere):

She give her the answer. This is kinda taking a long time

OpenStudy (tiniloisx):

I don't get it sorry :/ that last part

OpenStudy (tiniloisx):

Are we dividing rather than multiplying?

OpenStudy (tiniloisx):

@Nnesha @kidrah69 @quickstudent

OpenStudy (tiniloisx):

My theory: is 130/6 = $21.67 = how much she makes each hour $399 - $130 = 269 how much she has left to make I don't know where to go from here or if I am even correct.

OpenStudy (phi):

sometimes it is easier to figure out a simple problem first. For example, say you made $2 an hour and you want to make $4. How many hours do you have to work ?

OpenStudy (tiniloisx):

2 hours?

OpenStudy (phi):

yes $2 for the first hour, 2$ for the 2nd hour if you want to make $6 , how many hours ?

OpenStudy (tiniloisx):

You would have to work an additional hour so 3 hours?

OpenStudy (phi):

yes. so here is what you know: if you make 2 an hour, and you want to earn 6, you work 3 hours notice 6 hours/2 = 3 if we keep doing these simple problems, it will always be total dollars/ rate = # of hours

OpenStudy (phi):

in your problem $399 is the total the rate is 21.67 we want the number of hours

OpenStudy (phi):

another way to think of the problem is rate * time = amount 21.67 * time= 399 if we divide both sides by 21.67 (this is algebra) time= 399/21.67

OpenStudy (tiniloisx):

Okay that makes so much sense so we would get: 18.41 At least that is what I got.

OpenStudy (phi):

another way to do the problem is use ratios: \[ \frac{6 \ hrs}{130}= \frac{x}{399} \] and solve for x \[ x= 399 \cdot \frac{6}{130} \]

OpenStudy (phi):

yes, that is what I get, about 18.4 hours (18 hours and 25 minutes roughly)

OpenStudy (tiniloisx):

Thank you so much for explaining and very well detailed to the point where I actually got it! Wish you were my math teacher lol. I hope you have a fantastic day! :]

OpenStudy (phi):

ty